OpenStudy (anonymous):

The equation of a circle is (x + 7)2 + (y + 2)2 = 49. Determine the length of the diameter. (step by step please so i can learn)

OpenStudy (mikurout):

Do you know the equation of circle?

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 @mikurout

OpenStudy (mikurout):

Here r stands for which term?

OpenStudy (anonymous):

radius? @mikurout

OpenStudy (mikurout):

Yes

OpenStudy (bradely):

radius r = 7 diameter,d = 2r =2(7) =14 Source:www.mathskey.com

OpenStudy (mikurout):

Now you know the value of radius.

OpenStudy (mikurout):

As r^2 = 49 so r = \[\sqrt{49}\] =7

OpenStudy (anonymous):

so if the formula says "(x-h)' what causes the minus to change to addition? @mikurout

OpenStudy (mikurout):

It is the position of center

OpenStudy (anonymous):

so for example if the centers coordinates are (6,7) would that be addition or subtraction

OpenStudy (anonymous):

@mikurout

OpenStudy (mikurout):

substraction

OpenStudy (anonymous):

so when the cordinates are positive you subtract and when theyre negative youu add? ok i get it now thanks @mikurout

OpenStudy (mikurout):

Yes

OpenStudy (mikurout):

It is just the distance formula , that we are using for finding out the distance between center and any point on the circle.